OpenStudy (anonymous):

Combine like terms. What is a simpler form of the expression? -3(-y-2)+5y

OpenStudy (cggurumanjunath):

3y+6+5y=(3+5)y+6

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Csan you tell me the steps please???

OpenStudy (anonymous):

Can

OpenStudy (anonymous):

-3 (-y-2)+5y Multiply what's in parenthesis. -3 times -y =3y -3 times - 2 = 6 3y + 6 + 5y NOW, just combine like terms. 3 + 5 = 8 EQUATION ; 8y + 6

OpenStudy (anonymous):

Thanks